OpenStudy (anonymous):

If y varies directly with x and y = -7 when x = -1, find x when y = -84.

OpenStudy (anonymous):

y=kx -7=k *-1 solve for k -84= k*x solve for x. Remember k should be a number cuzz you olved for it :D

OpenStudy (anonymous):

-7=k*-1 -7/-1=k k=7 so -84=7*x -84/7=x x=-12